# Association of the IceCube neutrinos with CAZ blazar light curves

### v1

@@ Acknowledgements.
PK was supported by Academy of Finland projects 346071 and 345899. EL was supported by Academy of Finland projects 317636, 320045, and 346071. TH acknowledges support from the Academy of Finland projects 317383, 320085, 345899, and 362571 and from the European Union ERC-2024-COG - PARTICLES - 101169986. KK acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352). IL was funded by the European Union ERC-2022-STG - BOOTES - 101076343. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW. The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant #12540303 (PI: M.J.Graham). This work used data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The ATLAS project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. This work has made use of data from the Joan Oró Telescope (TJO) of the Montsec Observatory (OdM), which is owned by the Catalan Government and operated by the Institute for Space Studies of Catalonia (IEEC). This work makes use of Matplotlib (Hunter 2007), NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), and Astropy (Astropy Collaboration et al. 2022).References
@@
- Azzollini et al. (2025) Azzollini, A., Buson, S., Coleiro, A., et al. 2025, arXiv e-prints, arXiv:2507.03613
@@
- Kouch et al. (2025a) Kouch, P. M., Lindfors, E., Hovatta, T., & et al. 2025a, Submitted to A&A
- Kouch et al. (2025b) Kouch, P. M., Lindfors, E., Hovatta, T., et al. 2025b, A&A, 696, A73
@@
- Kovalev et al. (2025) Kovalev, Y. Y., Pushkarev, A. B., Gomez, J. L., et al. 2025, arXiv e-prints, arXiv:2504.09287
@@
- Lu et al. (2024) Lu, M.-X., Liang, Y.-F., Ouyang, X., Li, R.-L., & Wang, X.-G. 2024, arXiv e-prints, arXiv:2404.19730
@@
- Moretti & Caccianiga (2025) Moretti, A. & Caccianiga, A. 2025, arXiv e-prints, arXiv:2507.17908
@@
- Paggi et al. (2025) Paggi, A., Kouch, P. M., Liodakis, I., & et al. 2025, Submitted to A&A
- Paraschos et al. (2025) Paraschos, G. F., Traianou, E., Debbrecht, L. C., Liodakis, I., & Ros, E. 2025, arXiv e-prints, arXiv:2507.16929
@@
- Plavin et al. (2025) Plavin, A. V., Kovalev, Y. Y., & Troitsky, S. V. 2025, arXiv e-prints, arXiv:2503.08667
- Podlesnyi & Oikonomou (2025) Podlesnyi, E. & Oikonomou, F. 2025, arXiv e-prints, arXiv:2502.12111
@@
- Saurenhaus et al. (2025) Saurenhaus, L., Capel, F., Oikonomou, F., & Buchner, J. 2025, arXiv e-prints, arXiv:2507.06110
@@
- Stathopoulos & Petropoulou (2025) Stathopoulos, S. I. & Petropoulou, M. 2025, arXiv e-prints, arXiv:2507.08680
@@
Appendix A Low weight spatio-temporal associations
Here we describe and provide the light curve and sky map of three low-weight spatio-temporal associations arising from the BB95 at the peak of a BBHOP metric, similar to Sect. 4.5 where we did the same for high-weight spatio-temporal associations arising from the BB95 at the peak of a BBHOP metric.
The first of the low-weight spatio-temporal associations arising from the BB95 at the peak of a BBHOP metric is of the IBL CAZJ2227+0037 (4FGL J2227.9+0036) with IC180612A (=), whose light curve and sky map are given in Fig. 8. The neutrino IC180612A arrived at the peak of an outburst period. Intriguingly, CAZJ2227+0037 is associated with three other neutrinos: IC110807A, IC140114A, and IC200523A. These have weights of 0.266, 0.355, and 0.036, respectively. IC110807A arrived very close to the peak of another outburst period, but it unfortunately falls in a seasonal gap. This inadvertently excludes it from the spatio-temporal analysis as its temporal signal cannot be determined (for the adverse effects of seasonal gaps on our analysis, see Sect. 4.1). IC140114A also falls in a seasonal gap, although it likely did not arrive during a flare. Thus, this IBL has one definite, another likely, and another potential spatio-temporal association as well as one spatial association, which together make it an interesting neutrino-emitting candidate to follow up on. On top of this, its IBL nature fits the profile of the other spatio-temporally neutrino-associated sources well.

### v2

@@ Acknowledgements.
We thank the anonymous referee for constructive comments. PK was supported by Academy of Finland projects 346071 and 345899. EL was supported by Academy of Finland projects 317636, 320045, and 346071. TH acknowledges support from the Academy of Finland projects 317383, 320085, 345899, and 362571 and from the European Union ERC-2024-COG - PARTICLES - 101169986. KK acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 101002352). IL was funded by the European Union ERC-2022-STG - BOOTES - 101076343. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. Based on observations obtained with the Samuel Oschin Telescope 48-inch and the 60-inch Telescope at the Palomar Observatory as part of the Zwicky Transient Facility project. ZTF is supported by the National Science Foundation under Grant No. AST-2034437 and a collaboration including Caltech, IPAC, the Weizmann Institute for Science, the Oskar Klein Center at Stockholm University, the University of Maryland, Deutsches Elektronen-Synchrotron and Humboldt University, the TANGO Consortium of Taiwan, the University of Wisconsin at Milwaukee, Trinity College Dublin, Lawrence Livermore National Laboratories, and IN2P3, France. Operations are conducted by COO, IPAC, and UW. The ZTF forced-photometry service was funded under the Heising-Simons Foundation grant #12540303 (PI: M.J.Graham). This work used data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. The ATLAS project is primarily funded to search for near earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. This work was partially funded by Kepler/K2 grant J1944/80NSSC19K0112 and HST GO-15889, and STFC grants ST/T000198/1 and ST/S006109/1. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory, and The Millennium Institute of Astrophysics (MAS), Chile. This work has made use of data from the Joan Oró Telescope (TJO) of the Montsec Observatory (OdM), which is owned by the Catalan Government and operated by the Institute for Space Studies of Catalonia (IEEC). This work makes use of Matplotlib (Hunter 2007), NumPy (Harris et al. 2020), SciPy (Virtanen et al. 2020), and Astropy (Astropy Collaboration et al. 2022).References
@@
- Azzollini et al. (2025) Azzollini, A., Buson, S., Coleiro, A., et al. 2025, A&A, 700, A228
@@
- Huang et al. (2024) Huang, T. Q., Cao, Z., Chen, M., et al. 2024, in 38th International Cosmic Ray Conference, 1080
@@
- Kouch et al. (2025a) Kouch, P. M., Lindfors, E., Hovatta, T., et al. 2025a, A&A, 696, A73
@@
- Kouch et al. (2025b) Kouch, P. M., Lindfors, E., Hovatta, T., et al. 2025b, arXiv e-prints, arXiv:2510.16584
@@
- Kovalev et al. (2025) Kovalev, Y. Y., Pushkarev, A. B., Gómez, J. L., et al. 2025, A&A, 700, L12
@@
- Liodakis et al. (2019) Liodakis, I., Romani, R. W., Filippenko, A. V., Kocevski, D., & Zheng, W. 2019, ApJ, 880, 32
- Lu et al. (2025) Lu, M.-X., Liang, Y.-F., Ouyang, X., Li, R.-L., & Wang, X.-G. 2025, Phys. Rev. D, 112, 103013
@@
- Moretti & Caccianiga (2025) Moretti, A. & Caccianiga, A. 2025, A&A, 704, A184
@@
- Paraschos et al. (2025) Paraschos, G. F., Traianou, E., Debbrecht, L. C., Liodakis, I., & Ros, E. 2025, ApJ, 989, 208
@@
- Plavin et al. (2025) Plavin, A. V., Kovalev, Y. Y., & Troitsky, S. V. 2025, ApJ, 991, 33
- Podlesnyi & Oikonomou (2025) Podlesnyi, E. & Oikonomou, F. 2025, MNRAS, 544, 2897
@@
- Rodrigues et al. (2025) Rodrigues, X., Rieger, F., Bohdan, A., & Padovani, P. 2025, arXiv e-prints, arXiv:2508.18345
@@
- Saurenhaus et al. (2026) Saurenhaus, L., Capel, F., Oikonomou, F., & Buchner, J. 2026, Phys. Rev. D, 113, 023019
@@
- Stathopoulos & Petropoulou (2026) Stathopoulos, S. I. & Petropoulou, M. 2026, MNRAS, 545, staf2083
@@
- Zegarelli et al. (2025) Zegarelli, A., Franckowiak, A., Sommani, G., Valtonen-Mattila, N., & Yuan, T. 2025, arXiv e-prints, arXiv:2507.06176
@@
Appendix A Low weight spatiotemporal associations
Here, we give and describe the light curve and sky map of three low-weight spatiotemporal associations arising from the BB95 at the peak of a BBHOP metric, similar to Sect. 4.5 where we did the same for high-weight spatiotemporal associations arising from the BB95 at the peak of a BBHOP metric.
The first of the low-weight spatiotemporal associations arising from the BB95 at the peak of a BBHOP metric is of the IBL CAZJ2227+0037 (4FGL J2227.9+0036) with IC180612A (=), whose light curve and sky map are given in Fig. 8. The neutrino IC180612A arrived at the peak of an outburst period. Intriguingly, CAZJ2227+0037 is associated with three other neutrinos: IC110807A, IC140114A, and IC200523A. These have weights of 0.266, 0.355, and 0.036, respectively. IC110807A arrived very close to the peak of another outburst period, but it unfortunately falls in a seasonal gap. This inadvertently excludes it from the spatiotemporal analysis as its temporal signal cannot be determined (for the adverse effects of seasonal gaps on our analysis, see Sect. 4.1). IC140114A also falls in a seasonal gap, although it likely did not arrive during a flare. Thus, this IBL has one definite, another likely, and another potential spatiotemporal association as well as one spatial association, which together make it an interesting neutrino-emitting candidate to follow up on. On top of this, its IBL nature fits the profile of the other spatiotemporally neutrino-associated sources well.